# Hydrostatic Newton-Cartan Membranes

### v1

@@ Abstract
We construct the most general equilibrium partition function for Newton-Cartan membranes of codimension-1 defined on a flat ambient Newton-Cartan spacetime at second order in the hydrodynamic derivative expansion. We find the equilibrium constraints and the corresponding set of equilibrium constitutive relations. We show the explicit form of the Young-Laplace equation for a subset of non-vanishing and non-constant transport coefficients in the presence of a rotating fluid background with a constant vorticity.
I Introduction
Surfaces and interfaces are of importance on a wide range of physical phenomena such as biological systems Canham1970 ; Helfrich ; Tu2014 ; Cowin:2001 ; Mooendarbary:2013 ; Mow:1984 , condensed matter systems Kane2005 , and wave dynamics on spherical surfaces Streichan2017 ; Ritsma2014 ; Delplace2017 ; Shankar2017 ; Pearce2019 , such as planet Earth. A theoretical framework for studying such systems was established in Armas:2020 , based on the microscopic symmetry structure of the systems under study. A key aspect of the formalism used in Armas:2020 is that a geometric interpretation of temperature and mass chemical potential arise naturally from it. This allows for the study of elastic surfaces at finite temperature and mass chemical potential.
Following the work laid down in Armas:2020 , it is the goal of this paper to establish the equilibrium states of membranes that allow for exchange of both energy and matter. In particular, we will use a generating equilibrium partition function to write down the most general set of constitutive relations for membranes with finite temperature and chemical potential.
@@ I Introduction
The manuscript is organized as follows. In sections II and III we provide a summary of the necessary tools to understand the formalism presented in Armas:2020 . Section IV shows the use of the formalism to write down the equilibrium partition function up to second order in derivatives together with its corresponding equilibrium constraints and constitutive relations. In section V we write down the corresponding Young-Laplace equation in the presence of constant vorticity for a particular set of non-vanishing transport coefficient that depend on both temperature and chemical potential. Finally, in sectionVI we conclude with some comments about potential future work.
II Torsional Newton Cartan Background
@@
In this section we present an overview of Torsional Newton Cartan (TNC) geometry and introduce all the essential details that will be needed when using this geometry as the source of the mass, momentum, and energy currents. These details include a summary of TNC sub-manifolds, formalism that is needed to describe the equilibrium state of non-relativistic membranes. See Armas:2020 for a review [^1] Most of the overview presented in this and the next section, particularly regarding TNC sub-manifolds, can be considered a summary of what is discussed and introduced in Armas:2020 . of TNC geometry and TNC sub-manifolds.
II.1 TNC Geometry
@@
Let $\mathcal{M}_{d+1}$ be a (d+1)-dimensional manifold equipped with a TNC structure, consisting on the fields [^2] We use Greek indices to denote space-time indices $\{\mu,\nu,...\}=0,...,d$ $\left(\tau_{\mu},h_{\mu\nu},m_{\mu}\right)$; with $h_{\mu\nu}$ a symmetric spatial metric, $\tau_{\mu}$ a 1-form known as the clock 1-form, and $m_{\mu}$ a U(1) gauge connection associated to particle number conservation. An inverse structure $\left(\upsilon^{\mu},h^{\mu\nu}\right)$ can be defined following the orthonormality conditions
@@
The TNC structure on $\mathcal{M}_{d+1}$ in terms of the fields $\left(\tau_{\mu},h_{\mu\nu},m_{\mu}\right)$ transform under diffeomorphisms (local coordinate transformations parametrized by $\xi^{\mu}$), U(1) (mass gauge transformations parametrized by $\sigma$), Milne boosts (local rotations and local Galilean boosts parametrized by $\lambda_{\mu}$) as
@@
$e=\sqrt{\det\left(h_{\mu\nu}-\tau_{\mu}\tau_{\nu}\right)}.$ (6)
@@
To study dynamic evolution of systems coupled to a TNC geometry covariantly it is necessary to introduce a covariant derivative adapted to the generic curved TNC background. There exist no unique metric-compatible connection so we ask for compatibility with $\{\tau_{\mu},h^{\mu\nu}\}$, namely
@@
with $\nabla$ the covariant derivative constructed using the affine connection $\Gamma^{\rho}{}_{\mu\nu}$. A connection satisfying (16) and explicitely boost invariant is given by
@@
For most cases of physical interest such torsion vanishes [^3] This statement is equivalent to saying that for torsionless systems there exists a universal time.. The connection (17) is invariant under boosts, however it transforms under U(1) mass transformations as
@@
For a vanishing torsion the connection is invariant under both boost and U(1) transformations. The Riemann tensor $R^{\rho}{}_{\sigma\mu\nu}$ of the TNC background is defined in the standard way
@@
II.2 TNC Submanifolds
A (p+1)-dimensional TNC submanifold $\Sigma_{p+1}$, with coordinates [^4] We will use the lower case latin indices $\{a,b,...=0,...,p\}$ to denote submanifold spacetime indices $\sigma^{a}$, embedded on a (d+1)-dimensional TNC manifold $\mathcal{M}_{d+1}$, with coordinates $x^{\mu}$, is parametrized by the map $X^{\mu}:\Sigma\rightarrow\mathcal{M},\mu=0,...,d$ relating the coordinates $\sigma^{a}$ on $\Sigma_{p+1}$ to the location $X^{\mu}(\sigma^{a})$ on $\mathcal{M}$. In other words, the embedding specifies the location of the surface on the ambient space $\mathcal{M}$.
Tangent vectors $u^{\mu}_{a}$ to the surface are defined from the embedding map as
@@ II.2 TNC Submanifolds
The normal 1-forms [^5] The upper case latin indices $\{I,J,...=1,...,d-p\}$ correspond to indices on the transverse directions to the surface. We will consider codimension 1 surfaces meaning there will only be a single transverse direction and no upper latin indices will be used beyond this section. $n^{I}_{\mu}$ will be implicitly defined through the relations
@@
where $I=1,...,d-p$. Notice that $n^{I}_{\mu}$ is not uniquely fixed by (23) as such relation is invariant under SO(d-p) transformations. The inverse tangent vector $u^{a}_{\mu}$ and inverse normal 1-form $n^{\mu}_{I}$ can be introduce through the orthonormality conditions
@@
On time-like submanifolds the normal vector $n^{\mu}_{I}$ satisfies $n^{\mu}_{I}\tau_{\mu}=0$, therefore we can explicitly set $n^{\mu I}=h^{\mu\nu}n^{I}_{\nu}$. We can use the tangent vectors and normal 1-forms, as well as their inverses, to project any tensor tangentially or orthogonally to the surface. In particular the relevant boost-invariant TNC structure induced on the surface will be given by the set of objects $\{\tau_{a},\bar{h}_{ab},\check{\Phi}\}$ and $\{\hat{\upsilon}^{a},h^{ab}\}$ related to the ambient TNC structures through the relations
@@
Note that the underlying U(1) transformation is the one associated to the gauge connection $\check{m}_{a}$, $\delta\check{m}_{a}=\partial_{a}\sigma$ , related to the bulk one via
$\check{m}_{a}$ $=u^{\mu}_{a}m_{\mu}-\frac{1}{2}\upsilon^{I}\upsilon_{I}\tau_{a}.$ (41)
@@
$K_{ab}{}^{I}=-u^{\mu}_{a}u^{\nu}_{b}\nabla_{(\mu}n^{I}_{\nu)}=n^{I}_{\mu}D_{a}u^{\mu}_{b}+\frac{1}{2}\hat{\upsilon}^{I}\tau_{ab},$ (49)
and transforming under mass U(1) transformations as
@@
Notice that the covariant derivative $D$ does not act on normal indices $I$. It is also possible to define a covariant derivative $\mathcal{D}$ that acts on all types of indices. In particular we will find that
@@
The tensor $\omega_{a}{}^{I}{}_{J}$ is called the external rotation tensor and can be interpreted as a SO(d-p) connection, it is defined from the covariant derivative of the normal 1-form
@@
Notice that for submanifolds with codimension one the external rotation tensor vanishes by definition. The curvature associated to the external rotation tensor is known as the outer curvature $\Omega^{I}{}_{Jab}$ and is given by
@@
III Currents and conservation equations
@@
$S=S_{\Sigma}\left[\hat{\phi},\bar{h}_{ab},\tau_{a},K_{ab}{}^{I}\right],$ (59)
where $\phi$ are the fields describing the microscopic theory with $\hat{\phi}$ their value at the sub-manifold. A variation of this action will take the form [^6] Notice that there is no transverse index on $\mathcal{D}$ or $\delta K$ as the boundary is a codimension 1 surface.
@@
where $\hat{e}=\sqrt{\frac{\det\bar{h}_{ab}}{2\hat{\Phi}}}$, $\mathcal{T}^{a}_{S}$ is the surface boundary current, $\mathcal{T}^{ab}_{S}$ is the surface Cauchy stress-mass tensorGeracie:2016 , and $\mathcal{D}^{ab}$ is the bending moment encoding elastic responses. The equations of motion can be derived by performing a Lagrangian variation of the actionArmas:2020 , such variation transforms the bulk fields and keeps the embedding fixed. Namely, the Lagrangian variations under diffeomorphisms and U(1) mass transformations take the explicit form
@@
$\delta\bar{h}_{ab}$ $=2\bar{h}_{\rho(a}D_{b)}\xi^{\rho}-2\xi^{\sigma}\left[\bar{\mathcal{K}}_{\sigma(a}\tau_{b)}+\tau_{\sigma}\tau_{(a}D_{b)}\Phi+\tau_{a}\tau_{b}\nabla_{\sigma}\Phi+F_{\sigma(a}\tau_{b)}\Phi\right]-2\tau_{(a}\partial_{b)}\sigma\,,$ (62)
$\delta K_{ab}$ $=n_{\alpha}D_{a}D_{b}\xi^{\alpha}-R^{\alpha}{}_{ba\lambda}\xi^{\lambda}n_{\alpha}+\left(F_{\lambda b}n_{\beta}D_{a}\hat{\upsilon}^{\beta}+vu^{\beta}_{b}D_{a}F_{\beta\lambda}+vF_{\mu\lambda}n^{\mu}K_{ab}\right)\xi^{\lambda}$ (63)
$\hphantom{=}+vF_{b\alpha}D_{a}\xi^{\alpha}-\frac{1}{2}vF_{ab}n^{\alpha}n_{\lambda}D_{\alpha}\xi^{\lambda},$
where $\bar{\mathcal{K}}_{\mu\nu}=-\frac{1}{2}\mathcal{L}_{\hat{\upsilon}}\bar{h}_{\mu\nu}$. Invariance of (60) under this set of transformations gives out the derivative constraint
$F_{ab}=0\,,$ (64)
the tangential equations of motion
$D_{a}\left[\left(\mathcal{T}_{S}\right)^{ab}\tau_{a}\right]$ $=0\,,$ (65)
$D_{a}\left(\mathcal{T}_{S}\right)^{a}-\mathcal{K}^{\Sigma}_{ab}\left(\mathcal{T}_{S}\right)^{ab}-2\left(\mathcal{T}_{S}\right)^{ab}\tau_{b}D_{a}\hat{\Phi}$ $=2\hat{\upsilon}^{d}D_{a}\left(\mathcal{D}^{ab}K_{db}\right)-\hat{\upsilon}^{a}\mathcal{D}^{cd}D_{a}K_{cd},$ (66)
$D_{a}\left(\mathcal{T}_{S}\right)^{ac}$ $=h^{ca}\mathcal{D}^{cd}D_{a}K_{cd}-2h^{cd}D_{a}\left(\mathcal{D}^{ab}K_{db}\right),$ (67)
@@
$K_{ab}\left(\mathcal{T}_{S}\right)^{ab}$ $=D_{a}D_{b}\mathcal{D}^{ab}-h^{cd}\mathcal{D}^{ab}K_{ac}K_{bd}+\mathcal{D}^{ab}R_{ab}$ (68)
$+f_{a}\left[\left(\mathcal{T}_{S}\right)^{a}+vD_{b}\mathcal{D}^{ab}+\hat{\upsilon}^{b}K_{bd}\mathcal{D}^{ad}+2\mathcal{D}^{ab}D_{b}v\right],$
where we defined $\mathcal{K}^{\Sigma}_{ab}=-\frac{1}{2}\mathcal{L}_{\hat{\upsilon}}\bar{h}_{ab},R_{ab}=R^{\lambda}{}_{a\mu b}n^{\mu}n_{\lambda},$ and $f_{a}=F_{a\lambda}n^{\lambda}$. This last equation is the shape equation and will determine the embedding function $n_{\mu}X^{\mu}$.
IV Equilibrium Constituve Relations
Our focus will be on writing out the membrane constitutive relations on equilibrium configurations. Such configurations are characterized by the existence of a time-like killing vector $V^{\mu}$ that characterizes the time independence of the fluid configuration. In particular to have an equilibrium configuration the background should be invariant under the symmetry transformation

### v2

@@ Abstract
An application of the Newton-Cartan framework to the study of membranes is presented. Specifically, for membranes of co-dimension one in hydrostatic equilibrium embedded in a flat ambient Newton-Cartan spacetime. For such membranes, the corresponding equilibrium partition function at second order in the hydrodynamic derivative expansion is shown. Equilibrium constraints and the corresponding set of equilibrium constitutive relations are found. For the generically non-constant elastic subset of thermodynamic coefficients, the Young-Laplace equation is presented for the case of two-dimensional axisymmetric closed membranes embedded in a flat three-dimensional spacetime with constant ambient vorticity. Some numerical solutions to this Young-Laplace equation are examined, and some analytic solutions for particular choices of the thermodynamic coefficients are also discussed.
1 Introduction
Surfaces and interfaces are of importance in a wide range of physical phenomena such as biological systems [7, 17, 29, 10, 21, 22], condensed matter systems [20], and wave dynamics on spherical surfaces [28, 25, 11, 27, 24], such as planet Earth. Closed surfaces modeling physical shells and membranes in equilibrium can adopt different shapes depending on their free energy minimum. Typically, elastic contributions to the free energy will lead to spheres, tori, and biconcave discs, similar to the shape observed in red blood cells.
A theoretical framework for studying such systems in equilibrium was established in 2020 by Armas et al. [3]. The membrane is treated as a submanifold in a Newton-Cartan spacetime. A key aspect of the formalism is that a geometric interpretation of temperature and mass chemical potential arises naturally from it, which allows for the study of elastic surfaces at finite temperature and mass chemical potential. One of their results is a generalization of the Canham-Helfrich free energy for lipid bilayer membranes, which predicts the shapes mentioned above.
In the present manuscript, we make extensive use of this framework to derive general constitutive relations for fluid membranes of co-dimension one in hydrostatic equilibrium embedded in a flat ambient Newton-Cartan spacetime up to second order in the hydrodynamic derivative expansion. Hydrostatic conditions force the temperature to remain constant while the chemical potential is determined by the rotational kinetic energy of the fluid elements. The corresponding free-energy and constitutive relations can be derived, and the variation of such free-energy will lead to the Young-Laplace equation, which will determine the shape of the membrane. The Young-Laplace equation with purely elastic contributions is examined for two-dimensional axisymmetric membranes embedded in a spacetime with constant ambient vorticity. This corresponds to a fluid membrane immersed in a fluid that rotates with constant vorticity. Limiting the constitutive relations to the purely elastic subset results in the Young-Laplace equation depending on the usual elastic properties given by the surface tension, the bending and Gauss moduli, and the spontaneous curvature [23], but with non-trivial dependence on the chemical potential, which itself will depend on the vorticity through the equilibrium conditions. New thermodynamic parameters arise through these contributions, the effects of which have not been studied before, although the possibility of including them has been discussed before [3, 4]. Alternatively, part of the present work can be considered as the non-relativistic limit of the framework developed by Armas in [5] and [4].
This work contributes to the active research on membrane dynamics and demonstrates a simple application of the Newton-Cartan submanifold formalism, showing that hydrostatic membranes can sustain rotation and giving some examples of the shapes they can take in such flows. We find it worth mentioning the work by Salbreux and Jülicher, who derived, also from a geometrical standpoint, a covariant theory of active surfaces that appear in biological and chemical processes [26]. Their work provides a framework for studying membranes with microstructures that break symmetries, such as chiral surfaces. Symmetry considerations lead them to specific constitutive relations for stress and moment tensors. For liquid crystal membranes, Vafa and Mahadevan have proposed a framework to study the deformations produced by topological defects [30]. Their work can potentially model epithelial morphogenesis. The numerical studies of Hoffman, Carenza and Giomi on liquid crystal shells show that intense activity can lead to topological changes [18]. In particular, extensile activity can transform a flattened sphere into a torus, similarly to the effect of vorticity that was found in some of the numerical solutions shown in this work.
@@ 1 Introduction
The manuscript is organized as follows. In sections 2 and 3, a summary of the necessary tools to understand the formalism presented in [3] is provided. Section 4 shows the use of the formalism to write down the equilibrium partition function up to second order in derivatives together with its corresponding equilibrium constraints and constitutive relations. In section 5, we show the corresponding Young-Laplace equation in the presence of constant ambient vorticity for the elastic subset of thermodynamic coefficients assuming a dependence on both temperature and chemical potential. Finally, in section 6 we conclude with some comments on potential future work.
2 Torsional Newton Cartan Background
@@
In this section, we present an overview of Torsional Newton Cartan (TNC) geometry  [^1] Newton-Cartan geometry was first introduced by Cartan in [8, 9]. A modern understanding of non-relativistic gravity and non-relativistic geometry was pioneered in [2], with more systematic approaches done in works such as [31, 13, 14]. See [15] for a modern review of Newton-Cartan geometry.and introduce all the essential details that will be needed when using this geometry as the source of the mass, momentum, and energy currents. The essential details include a summary of TNC sub-manifolds, a formalism that is needed to describe the equilibrium state of non-relativistic membranes. The overview presented in this and the next section regarding TNC sub-manifolds can be considered a summary of the formalism of Newton-Cartan submanifolds that was introduced in [3].
2.1 TNC Geometry
@@
Let $\mathcal{M}_{d+1}$ be a (d+1)-dimensional manifold equipped with a TNC structure, consisting on the fields [^2] We use Greek indices to denote space-time indices $\{\mu,\nu,...\}=0,...,d$ $\left(\tau_{\mu},h_{\mu\nu},m_{\mu}\right)$; with $h_{\mu\nu}$ a symmetric spatial metric, $\tau_{\mu}$ a 1-form known as the clock 1-form, and $m_{\mu}$ a U(1) gauge connection associated with particle number conservation. An inverse structure $\left(\upsilon^{\mu},h^{\mu\nu}\right)$ can be defined following the orthonormality conditions
@@
The TNC structure on $\mathcal{M}_{d+1}$ in terms of the fields $\left(\tau_{\mu},h_{\mu\nu},m_{\mu}\right)$ transform under diffeomorphisms (local coordinate transformations parametrized by $\xi^{\mu}$), a U(1) transformation (mass gauge transformation parametrized by $\sigma$), and Milne boosts (local Galilean boosts parametrized by $\lambda_{\mu}$) as
@@
$e=\sqrt{-\det\left(h_{\mu\nu}-\tau_{\mu}\tau_{\nu}\right)}.$ (6)
@@
To study the dynamic evolution of systems coupled to a TNC geometry covariantly, it is necessary to introduce a covariant derivative adapted to the generic curved TNC background. There exists no unique metric-compatible connection [6, 16], so we ask for compatibility with $\{\tau_{\mu},h^{\mu\nu}\}$, namely
@@
with $\nabla$ the covariant derivative constructed using the affine connection $\Gamma^{\rho}{}_{\mu\nu}$. A connection satisfying (16) and explicitly boost invariant is given by
@@
For most cases of physical interest, such torsion vanishes  [^3] This statement is equivalent to saying that for torsionless systems there exists a universal time.. The connection (17) is invariant under boosts, however it transforms under U(1) mass transformations as
@@
For a vanishing torsion, the connection is invariant under both boost and U(1) transformations. The Riemann tensor $R^{\rho}{}_{\sigma\mu\nu}$ of the TNC background is defined in the standard way
@@
2.2 TNC Submanifolds
A (p+1)-dimensional TNC submanifold $\Sigma_{p+1}$, with coordinates  [^4] We will use the lower case latin indices $\{a,b,...=0,...,p\}$ to denote the submanifold spacetime indices $\sigma^{a}$, embedded in a (d+1)-dimensional TNC manifold $\mathcal{M}_{d+1}$, with coordinates $x^{\mu}$, is parametrized by the map $X^{\mu}:\Sigma\rightarrow\mathcal{M},\mu=0,...,d$ relating the coordinates $\sigma^{a}$ on $\Sigma_{p+1}$ to the location $X^{\mu}(\sigma^{a})$ on $\mathcal{M}$. In other words, the embedding specifies the location of the surface in the ambient space $\mathcal{M}$. The tangent vectors $u^{\mu}_{a}$ to the surface are defined from the embedding map as
@@ 2.2 TNC Submanifolds
The normal 1-forms [^5] The upper case Latin indices $\{I,J,...=1,...,d-p\}$ correspond to indices on the transverse directions to the surface. We will consider co-dimension one surfaces, meaning there will only be a single transverse direction, and no upper Latin indices will be used beyond this section. $n^{I}_{\mu}$ will be implicitly defined through the relations
@@
where I=1,,dpI=1,...,d-p. Notice that $n^{I}_{\mu}$ is not uniquely fixed by (23) as such relation is invariant under SO(d-p) transformations. The inverse tangent vector $u^{a}_{\mu}$ and inverse normal 1-form $n^{\mu}_{I}$ can be introduced through the orthonormality conditions
@@
In time-like submanifolds the normal vector $n^{\mu}_{I}$ satisfies $n^{\mu}_{I}\tau_{\mu}=0$, therefore we can explicitly set $n^{\mu I}=h^{\mu\nu}n^{I}_{\nu}$. We can use the tangent vectors and normal 1-forms, as well as their inverses, to project any tensor tangentially or orthogonally to the surface. In particular, the relevant boost-invariant TNC structure induced on the surface will be given by the set of objects $\{\tau_{a},\bar{h}_{ab},\check{\Phi}\}$ and $\{\hat{\upsilon}^{a},h^{ab}\}$ related to the ambient TNC structures through the relations
@@
Note that the underlying U(1) transformation is the one associated with the gauge connection $\check{m}_{a}$, $\delta\check{m}_{a}=\partial_{a}\sigma$, related to the bulk one via
$\check{m}_{a}$ $=u^{\mu}_{a}m_{\mu}-\frac{1}{2}\upsilon^{I}\upsilon_{I}\tau_{a}.\$(41)
@@
$K_{ab}{}^{I}=-u^{\mu}_{a}u^{\nu}_{b}\nabla_{(\mu}n^{I}_{\nu)}=n^{I}_{\mu}D_{a}u^{\mu}_{b}+\frac{1}{2}\hat{\upsilon}^{I}F_{ab},$ (49)
that transforms under mass U(1) transformations as
@@
Notice that the covariant derivative $D$ does not act on normal indices $I$. It is also possible to define a covariant derivative $\mathcal{D}$ that acts on all types of indices. In particular, it is found that
@@
The tensor $\omega_{a}{}^{I}{}_{J}$ is called the external rotation tensor and can be interpreted as an SO(d-p) connection. It is defined from the covariant derivative of the normal 1-form
@@
Notice that for submanifolds with co-dimension one the external rotation tensor vanishes by definition. The curvature associated with the external rotation tensor is known as the outer curvature $\Omega^{I}{}_{Jab}$ and is given by
@@
3 Currents and conservation equations
@@
$S=S_{\Sigma}\left[\hat{\phi},\bar{h}_{ab},\tau_{a},K_{ab}\right],$ (59)
where $\phi$ are the fields describing the microscopic theory with $\hat{\phi}$ their value at the sub-manifold. A variation of this action will take the form [^6] Notice that there is no transverse index on $\mathcal{D}$ or $\delta K$ as the boundary is a co-dimension one surface.
@@
where $\hat{e}=\sqrt{\frac{\det\bar{h}_{ab}}{2\hat{\Phi}}}$, $\mathcal{T}^{a}_{S}$ is the surface boundary current, $\mathcal{T}^{ab}_{S}$ is the surface Cauchy stress-mass tensor[12], and $\mathcal{D}^{ab}$ is the bending moment encoding elastic responses. he conservation equation associated with mass conservation can be derived by looking at the U(1) mass transformations
$\delta\bar{h}_{ab}$ $=-\tau_{a}\partial_{b}\sigma-\tau_{b}\partial_{a}\sigma\,$, (61)
$\delta K_{ab}$ $=-\frac{1}{2}f_{a}\partial_{b}\sigma-\frac{1}{2}f_{b}\partial_{a}\sigma\,$, (62)
with $f_{a}=F_{a\lambda}n^{\lambda}$, and asking for (60) to be invariant under this set of transformations. This results in the mass conservation equation
$D_{a}\left[\left(\mathcal{T}_{S}\right)^{ab}\tau_{b}+\mathcal{D}^{ab}f_{b}\right]=0\,.$ (63)
The remaining equations of motion can be derived by considering Lagrangian variations [5, 3] in (60). Lagrangian variations are those that transform the bulk fields but keep the embedding fixed. The Lagrangian variations under diffeomorphisms relevant to (60) are
@@
$\delta\bar{h}_{ab}$ $=2\bar{h}_{\rho(a}D_{b)}\xi^{\rho}-2\xi^{\sigma}\left[\bar{\mathcal{K}}_{\sigma(a}\tau_{b)}+\tau_{\sigma}\tau_{(a}D_{b)}\Phi\right$. (65)
$\left.+\tau_{a}\tau_{b}\nabla_{\sigma}\Phi+F_{\sigma(a}\tau_{b)}\Phi\right]\,$,
$\delta K_{ab}$ $=n_{\mu}D_{a}D_{b}\xi^{\mu}-R^{\mu}{}_{ba\alpha}\xi^{\alpha}n_{\mu}-vK_{ab}F_{\mu\lambda}\xi^{\mu}n^{\lambda}$ (66)
$+n_{\mu}u^{\lambda}_{b}D_{a}\left(\hat{\upsilon}^{\mu}F_{\lambda\alpha}\xi^{\alpha}\right)+\frac{1}{2}vn_{\mu}F_{ab}n^{\lambda}\nabla_{\lambda}\xi^{\mu}\,$,
where $\bar{\mathcal{K}}_{\mu\nu}=-\frac{1}{2}\mathcal{L}_{\hat{\upsilon}}\bar{h}_{\mu\nu}$ and $v=\hat{\upsilon}^{m}n_{\mu}$. Invariance of (60) under this set of transformations gives the tangential equations of motion
$D_{a}\left(\mathcal{T}_{S}\right)^{a}$ $=\mathcal{K}^{\Sigma}_{ab}\left(\mathcal{T}_{S}\right)^{ab}+2\left(\mathcal{T}_{S}\right)^{ab}\tau_{b}D_{a}\check{\Phi}$ (67)
$+2\hat{\upsilon}^{d}D_{a}\left(\mathcal{D}^{ab}K_{db}\right)-\hat{\upsilon}^{a}\mathcal{D}^{cd}D_{a}K_{cd}$
$-2\check{\Phi}D_{a}\left(\mathcal{D}^{ab}f_{b}\right)-2\mathfrak{a}_{c}\left(\mathcal{T}_{S}\right)^{c}+2\check{\Phi}\mathfrak{a}_{c}\tau_{d}\left(\mathcal{T}_{S}\right)^{cd}$
$+2\mathfrak{a}_{c}\hat{\upsilon}^{a}K_{ad}\mathcal{D}^{cd}\,$,
$D_{a}\left(\mathcal{T}_{S}\right)^{ac}$ $=h^{ca}\mathcal{D}^{fg}D_{a}K_{fg}-2h^{cd}D_{a}\left(\mathcal{D}^{ab}K_{db}\right)$ (68)
$-\mathfrak{a}_{f}h^{cb}\bar{h}_{bd}\left(\mathcal{T}_{S}\right)^{df}-F_{ab}h^{bc}\left(\mathcal{T}_{S}\right)^{a}+\hat{\upsilon}^{c}D_{a}\left(\mathcal{D}^{ab}f_{b}\right)$
$-2F_{db}h^{cd}\hat{\upsilon}^{f}K_{fa}\mathcal{D}^{ab}\,$,
@@
$K_{ab}\left(\mathcal{T}_{S}\right)^{ab}$ $=D_{a}D_{b}\mathcal{D}^{ab}-h^{cd}\mathcal{D}^{ab}K_{ac}K_{bd}$ (69)
$+\mathcal{D}^{ab}R_{ab}+vD_{a}\left(\mathcal{D}^{ab}f_{b}\right)-v\mathfrak{a}_{c}\left(\mathcal{T}_{S}\right)^{cd}\tau_{d}$
$-f_{a}\left[\left(\mathcal{T}_{S}\right)^{a}+vD_{b}\mathcal{D}^{ab}-\hat{\upsilon}^{b}K_{bd}\mathcal{D}^{ad}-\frac{v}{2}\mathfrak{a}_{b}\mathcal{D}^{ab}\right]$
$-vF_{ab}h^{ac}\left(K_{cd}+\frac{v}{4}F_{cd}\right)\mathcal{D}^{bd}\,$,
where we defined $\mathcal{K}^{\Sigma}_{ab}=-\frac{1}{2}\mathcal{L}_{\hat{\upsilon}}\bar{h}_{ab},R_{ab}=R^{\lambda}{}_{a\mu b}n^{\mu}n_{\lambda},$ and $\mathfrak{a}_{n}=\hat{\upsilon}^{m}F_{mn}$. This last equation is the shape equation and will determine the embedding function $n_{\mu}X^{\mu}$.
4 Equilibrium Constitutive Relations
Our focus will be on writing out the membrane constitutive relations on equilibrium configurations. Such configurations are characterized by the existence of a time-like killing vector $V^{\mu}$ that characterizes the time independence of the fluid configuration. In particular, to have an equilibrium configuration, the background should be invariant under the symmetry transformation